# Automatically Labeling \$200B Life-Saving Datasets: A Large Clinical Trial Outcome Benchmark

### v2

@@ Abstract
Background The global cost of drug discovery and development exceeds \$200 billion annually, with clinical trial outcomes playing a critical role in the regulatory approval of new drugs and impacting patient outcomes. Despite their significance, large-scale, high-quality clinical trial outcome data are not readily available to the public, limiting advances in trial outcome predictive modeling.
Methods We introduce the Clinical Trial Outcome (CTO) knowledge base, a fully reproducible, large-scale (around 125K drug and biologics trials), open-source of clinical trial information including large language model (LLM) interpretations of publications, matched trials over phases, sentiment analysis from news, stock prices of trial sponsors, and other trial-related metrics. From this knowledge base, we additionally performed manual annotation of a set of recent clinical trials from 2020-2024.
Results We evaluated the quality of our knowledge base by generating high-quality trial outcome labels that demonstrate strong agreement with previously published expert annotations, achieving an F1 score of 94 for Phase 3 trials and 91 across all phases. Additionally, we benchmarked a suite of standard machine learning models on our manually annotated set, highlighting the distribution shift of recent trials and the need for continuously updated labeling methods.
Conclusions By analyzing CTO’s performance on recent trials, we showed a need for recent, high-quality trial outcome labels. We release our knowledge base and labels to the public at https://chufangao.github.io/CTOD, which will also be regularly updated to support ongoing research in clinical trial outcomes, offering insights that could optimize the drug development process.
@@ Introduction
A clinical trial is an indispensable step toward developing a new drug, involving human participants to test the drug’s efficacy and safety for treating target diseases. In 2022, drug discovery and development spending reached 244 billion dollars globally [1]. Among these, the clinical trial market reached \$44.3 billion in 2020 and is expected to grow to \$69.3 billion by 2028 [2]. Low efficacy, safety issues, and poor trial protocol design can lead to trial failures [3, 4, 5]. Eroom’s Law111reverse of “Moore’s Law” shows that the number of new FDA-approved drugs per billion US dollars of R&D spending has halved approximately every nine years since 1950, even with inflation adjustment [6]. Despite this decline in approval efficiency, the volume of drug and biologics trials continues to grow each year. However, obtaining high-quality, human-labeled trial data remains expensive and largely inaccessible due to the limited public availability of labels. Given these challenges, predicting trial outcomes in silico—using computational methods—could significantly enhance drug development efficiency.
@@
The FDA does not release the clinical trial ID (NCTID) in documents of approved drug applications. Other public data sources, such as the ClinicalTrials.gov database with more than 500,000 historical trials [7, 8, 9], provide vital information for identifying trial recruitment and design information but does not provide a concrete label for trial outcome directly. The absence of a centralized, easily accessible knowledge base that consolidates clinical trial outcomes, drug approvals, and intervention phases poses a significant challenge for researchers and drug developers [10]. This fragmented information landscape hinders the development of accurate predictive models and can slow down the drug discovery and development process. Furthermore, the number of clinical trials conducted each year continues to rise, with their data dynamically evolving under the influence of various external factors, such as technological innovations and global events like the COVID-19 pandemic. This ever-changing landscape underscores the need for updated trial outcome labels to develop robust and high-performing clinical trial outcome prediction models.
@@
Trial Outcome Definition: Clinical trial outcomes are multifaceted and have diverse implications. These outcomes can involve meeting the primary endpoint as defined in the study, advancing to the next phase of the trial, obtaining regulatory approval, impacting the financial outcome for the sponsor (either positively or negatively), and influencing patient outcomes such as adverse effects and trial dropouts. Our paper follows the previous conventions [11, 12, 3, 13, 14, 15] and defines the trial outcome as a binary indicator, showing whether the trial achieves its primary endpoints and can progress to the next stage of drug development. For Phase 1 and 2 trials, success may mean moving to the next phase, such as from Phase 1 to Phase 2, and from Phase 2 to Phase 3. In Phase 3, success is measured by regulatory approval.
@@
Recently, initial efforts have been made to forecast specific aspects of clinical trials to enhance their outcomes. These efforts include employing electroencephalographic (EEG) measures to predict the effects of antidepressant treatments [16], optimizing drug toxicity predictions based on drug and target properties [17], and using phase II trial results to anticipate phase III trial outcomes [18]. Additionally, there is a growing interest in developing comprehensive methods for predicting trial outcomes. For example, predicting drug approvals for 15 disease groups by analyzing drug and clinical trial features using classical machine learning techniques [14], using multimodal drug structure and text information to predict outcomes based on a supervised set of data [11], an algorithm that computes the probability of technical success via asking experts a standardized questionnaire [19], and multimodal trial outcome prediction via omics, text, clinical trial design, and small molecule properties [15].
@@
To address these limitations, we propose CTO, the first large-scale, publicly available, open-source, and fully reproducible dataset of clinical trial outcomes. We compiled a comprehensive trial knowledge base by integrating diverse data sources, including trial-related publications, news headlines, stock prices of trial sponsors, and other relevant trial metrics. From this knowledge base, we derive weakly supervised trial outcome labels using the CTO’s automated labeling framework, which aggregates labels from trial phase linkages, LLM interpretations on publications, sentiment analysis of news headlines, p-values, and trial status, among other indicators. We validate CTO ’s labeling accuracy against previously published trial outcomes, achieving an F1 score of 91 across all trial phases, closely aligned with human-annotated clinical trial outcomes. Additionally, aided by the CTO’s framework, we manually curated a set of approximately 10,000 recent trials completed between 2020 and 2024. To ensure reproducibility and keep pace with rapidly evolving trial data, we share monthly updates of our trial knowledge base and labels generated by CTO ’s automated labeling framework.
Figure 1: Overview of our CTO benchmark pipeline. (A) Trial selection criteria used to curate a targeted dataset from the CTTI, narrowing from an initial 513.4K trials to 125.8K trials in CTO to a final manually curated set of 11K trials. (B) Trial knowledge base creation pipeline, including extraction of linked PubMed abstracts, recent abstract searches, trial metrics, news articles, and stock prices. (C) Clinical trial outcome labeling process, integrating the CTO automated labeling framework assisted manual labeling to generate curated gold standard labels. (D) The manual curation process, our rule-based annotation criteria, followed by manual checking on challenging clinical trials.
@@ Methods
In our study, we focused primarily on clinical trials involving drugs and biologics. Starting with a dataset of approximately 513,485 clinical trials from the CTTI, we applied a series of selection criteria to create a refined subset as illustrated in Figure 1A. First, we selected only trials with interventions categorized as drugs or biologics, reducing the dataset to 209,124 trials. Next, we excluded active trials and those still recruiting to focus exclusively on completed studies, resulting in 145,464 trials. Subsequently, trials without phase information were removed, further narrowing the dataset to 125,840 trials, which represents the number of trials labeled using our CTO’s automated labeling framework. To construct a manually curated test set, we then selected trials completed after 2020, yielding a subset of approximately 30,900 recent trials. Importantly, we ensured that this set had no overlap with the TOP benchmark. Finally, we performed a sampling across different trial phases and conditions, producing a manually curated set of 11,012 trials.
@@
Based on the CTTI data, we generate a trial knowledge base by aggregating and linking various data sources to the trials, as shown in Figure 1B. Our pipeline enables regular updates of the knowledge base given the updated CTTI data in a monthly frequency. The trial knowledge base extraction pipeline consists of four components: 1) PubMed abstracts extraction, 2) News extractions, 3) Trial metrics and 4) Stock price information extraction.
Pubmed abstract extraction: PubMed abstracts have been automatically linked to trials by the Clinical Trials Transformation Initiative (CTTI)[20, 21, 22] as well as other efforts [23]. Despite these efforts, we found that many recent abstracts were not linked to their corresponding trials. First, we extracted all linked PubMed abstracts for each trial through the NCBI API  [^2] https://ncbi.nlm.nih.gov/, and the statistics of the linked abstracts are given in the Supplementary. These abstracts can be categorized into 1) Background, 2) Derived, and 3) Results. Since we are interested in clinical trial outcomes, we only utilized abstracts in the Derived and Results categories. To further enrich the knowledge base, we performed regular searches on NCBI using NCT IDs to find newly published abstracts, linking any relevant ones to their associated trials. The initial count of drug or biologic trials with extracted PubMed abstracts linked from CTTI data was 49,277. Our additional search and extraction process linked 10,232 new trials with their recent PubMed abstracts. Ultimately, our knowledge base comprises 41,228 Results-type abstracts, 65,327 Derived-type abstracts, and 20,512 newly linked abstracts.
News extractions: In our initial analysis, we identified news sources as a valuable resource for determining the outcome of clinical trials, especially since the majority of trials in the CTTI dataset lack reported results. We extracted news articles related to completed drug and biologic trials by web scraping and linked them to the trials in our knowledge base. We sent requests to Google News for all completed drug and biologic trials (excluding phase 4), which accounted for around 85K trials. Due to rate limitations, we used an external API (SerpAPI) to streamline this process. Our findings revealed that most trials had no corresponding news coverage; only 3,688 trials were associated with any news articles.
Trial metrics: We extracted trial metrics from CTTI[24] as well as mining clinicaltrials.gov. Specifically, we extracted a total of 9 numerical metrics–number of patients, sponsors, sites, amendments, patient drops, deaths, serious adverse events, all adverse events, and the number of days the trial has been last updated after completion. We also extracted 3 binary metrics–whether results were reported, the completion/terminated status of the trial, and finally, whether the trial had a significant p-value. Detailed descriptions of each are in the supplementary.
@@
We developed a clinical trial labeling pipeline that first assigns rule-based labels. First, trials with statuses indicating ‘terminated’, ‘withdrawn’, ‘suspended’, ‘no longer available’} were labeled as failures. Trials labeled as ’approved for marketing’ were marked as successes. For trials with ‘completed’ as status, we examined the reported results: if any reported p-value for a primary endpoint was $<$0.05, the trial was labeled as a success. Unlabeled trials are then automatically labeled. The main idea behind our automated labeling framework is to combine multiple sources of labels, which have varying levels of accuracy. We call these sources of labels “weakly supervised labeling functions” or LFs for short[25]. Following this, we apply a label model that leverages our sources to refine labels and produce denoised pseudo-labels. Each LF, such as detecting efficacy based on p-value thresholds, assigns labels or abstains when uncertain. Phase-specific thresholds are also applied to certain metrics, such as adverse events, to account for variations in participant numbers across trial phases.
@@ Results
Table 1: Evaluation of automatically aggregated knowledge base data sources. Agreement of CTO with TOP (on TOP test split). $\text{CTO}_{MV}$ denotes Majority Vote, $\text{CTO}_{DP}$ denotes Data Programming [25], and $\text{CTO}_{RF}$ denotes Random Forest. Note that RF actively trained on the training human labels, whereas MV and DP are not. Bolded denotes best performance.
@@
I $\text{CTO}_{MV}$ 0.726 0.490
$\text{CTO}_{DP}$ 0.870 0.700
$\text{CTO}_{RF}$ 0.913 0.790
II $\text{CTO}_{MV}$ 0.689 0.430
$\text{CTO}_{DP}$ 0.856 0.623
$\text{CTO}_{RF}$ 0.878 0.693
III $\text{CTO}_{MV}$ 0.904 0.606
$\text{CTO}_{DP}$ 0.921 0.582
$\text{CTO}_{RF}$ 0.941 0.710
All $\text{CTO}_{MV}$ 0.793 0.529
$\text{CTO}_{DP}$ 0.884 0.646
$\text{CTO}_{RF}$ 0.909 0.729
We evaluate our knowledge base and labeling framework by comparing the agreement between our automatically aggregated labels against the expert-annotated TOP test set. Table 1 summarizes the performance of various label aggregation methods across different trial phases. The label aggregation methods evaluated include Majority Vote (CTOMV), Data Programming (CTODP), and Random Forest-based aggregation (CTORF). Across all trial phases, the Random Forest-based approach consistently outperforms the other methods, achieving the highest scores in both F1 and Cohen’s kappa metrics. CTORF achieves an F1 score of $0.909$ and a Cohen’s kappa ($\kappa$) of $0.729$ when evaluated across all phases, indicating substantial agreement [26]. Also, the phase-specific analysis further underscores the robustness of CTORF. It achieves F1 scores of $0.913$, $0.878$, and $0.941$ in phases I, II, and III, respectively. Additionally, the F1 score of 0.909 is much higher than any previous SOTA trial outcome prediction model [27, 12].
@@
Since automated labeling is never a replacement for humans, we manually annotated over 2,500 challenging cases using our knowledge base to create a high-confidence gold standard set for evaluation. Our process of creating this benchmark is shown in Figure 1D. We follow the same rule-based termination and p-value$<$0.05 as in the automated labeling step to remove the easy cases. For the rest of the unlabeled trials, we utilized 3 clinical trial experts to label the unlabeled trials by assessing related publications, news articles, and other reported outcomes to determine the trial outcome. In this benchmark, we evaluated the labels generated by the CTO automated labeling framework against human labels. The CTO framework demonstrated strong overall performance, achieving an F1 of 97.1. Notably, even in hard and challenging cases (after p-value and status filtering), where accurate labeling is more complex, CTO maintained an F1 of 72.7.
@@
Updated Trial Outcome Labels are Highly Valuable
@@
Figure 3: (A) Year-by-year comparison of trial outcome prediction performance for baseline models trained on the TOP dataset, the constantly updated CTO dataset, and our manually curated set, evaluated on trials completed from 2021 to 2024. The TOP dataset is fully utilized for training, whereas for the CTO and manually curated sets, only trials completed prior to each test year are used for training. (B) Top 10 Condition-wise performance of baselines on manual labels, with models trained on data $<$ 2022 and tested on $\geq$ 2022 using TOP, CTO, and manually curated set. (C) Similar to B, Condition-wise performance on the next 10 (Top 11-20 conditions)
@@
Given that we have manual labels for recent trials, we conducted experiments comparing the impact of our constantly updated CTO dataset against the previously publicly available fixed benchmark TOP dataset. We trained standard baseline trial outcome prediction models on the entire TOP dataset (train + valid + test) and assessed their year-by-year performance on trials completed between 2021 and 2024 in our manually curated dataset. For the CTO dataset, the baselines were trained using only trials completed prior to each test year, with training labels generated through our automated labeling framework. Similarly, we trained the baselines on the manually curated dataset using human-labeled outcomes for trials completed before each test year. The phase-wise performance comparison of these baseline models across the different training datasets is presented in Figure 3A.
@@
Baseline models trained using labels from the CTO dataset consistently outperform their counterparts trained on the TOP dataset across all years and trial phases. The CTO dataset significantly enhances model performance compared to the TOP dataset, underscoring the importance of constantly updated labels. This improvement is largely attributable to distribution shifts caused by evolving trial practices, therapeutic areas, and advancements in trial designs over time.
@@
This shows that CTO provides a scalable solution for labeling recent trial data, overcoming the infeasibility and time demands of obtaining new human labels. By achieving high agreement with human labels, CTO enables models to obtain close to human-level performance through leveraging large datasets to adapt to the evolving landscape of clinical trials.
Automated Labels for Condition-Specific Outcome Prediction
@@
Figure 3B and C present the performance of baseline models trained using our human-labeled data across the top 20 conditions in the dataset (3B shows conditions 1-10 and 3C shows conditions 11-20). Here, we train the baselines on trials completed before 2022 over TOP, CTO, and manually curated. The baselines are tested on the manually curated trials completed after 2022. The performance of the baseline models trained on CTO closely follows the models trained on our human labels. This demonstrates that the labels produced by CTO align closely with manual annotations despite being generated independently through an automated process. Moreover, these CTO-based models consistently outperform those trained on the TOP dataset, demonstrating the value of using a more current dataset. Furthermore, CTO facilitates the development of condition-specific clinical trial outcome prediction models by rapidly generating automated labels for condition-specific datasets, significantly reducing the time and effort involved in manual labeling, with CTO-trained baselines matching or even exceeding baselines trained on manually-curated in some cases.

### v3

# Automatically Labeling Clinical Trial Outcomes: A Large-Scale Benchmark for Drug Development

@@ Abstract
Background The high cost and complexity of drug discovery and development make clinical trial outcomes critical for regulatory approval and patient care. However, the limited availability of large-scale, high-quality clinical trial outcome data hinders advancements in predictive modeling and evidence-based decision-making.
Methods We introduce the Clinical Trial Outcome (CTO) benchmark, a fully reproducible, large-scale repository covering approximately 125,000 drug and biologics trials. CTO integrates large language model (LLM) interpretations of trial publications, phase progression tracking, sentiment analysis from news sources, stock price movements of trial sponsors, and additional trial-related metrics. To enhance outcome label quality and reliability, we manually annotated a dataset of clinical trials conducted between 2020 and 2024.
Results The CTO benchmark demonstrates high agreement with expert annotations, achieving an F1 score of 94 for Phase 3 trials and 91 across all phases. Additionally, benchmarking standard machine learning models on our manually annotated dataset reveals distribution shifts in recent trials, emphasizing the need for continuous updates in trial outcome labeling.
Conclusions Our analysis of CTO’s performance on recent clinical trials highlights the ongoing need for high-quality, up-to-date trial outcome labels. We publicly release the CTO knowledge base and annotated labels at https://chufangao.github.io/CTOD, with regular updates to support research on clinical trial outcomes and drive data-driven advancements in drug development.
@@ Introduction
Clinical trials are an indispensable yet high-risk stage in drug development, characterized by enormous costs and lengthy durations with only a small fraction of trials ultimately succeeding. Many trials fail due to issues of drug inefficacy, safety concerns, or challenges in patient recruitment. In 2022, drug discovery and development spending reached 244 billion dollars globally [1]. Among these, the clinical trial market reached \$44.3 billion in 2020 and is expected to grow to \$69.3 billion by 2028 [2]. Low efficacy, safety issues, and poor trial protocol design can lead to trial failures [3, 4, 5]. Eroom’s Law111reverse of “Moore’s Law” shows that the number of new FDA-approved drugs per billion US dollars of R&D spending has halved approximately every nine years since 1950, even with inflation adjustment [6]. Despite this decline in approval efficiency, the volume of drug and biologics trials continues to grow each year. However, obtaining high-quality, human-labeled trial data remains expensive and largely inaccessible due to the limited public availability of detailed trial description and their outcomes. Addressing these challenges, collecting diverse and comprehensive clinical trial datasets and using AI models to automatically assess trial outcomes could help establish a high-quality benchmark for clinical trials, ultimately enhancing the efficiency of drug development.
@@
The FDA does not release the clinical trial ID (NCTID) in documents of approved drug applications. Other public data sources, such as the ClinicalTrials.gov database with more than 500,000 historical trials [7, 8, 9], provide vital information for identifying trial recruitment and design information but does not provide comprehensive trial outcome directly [^2] The status is classified into basic categories: recruiting, completed, and terminated, but it does not clearly indicate whether the trial achieved the desired endpoints.. The absence of a centralized, easily accessible knowledge base that consolidates clinical trial outcomes, drug approvals, and trial phase transitions poses a significant challenge for researchers and drug developers [10]. This fragmented information landscape hinders the development of accurate predictive models and can slow down the drug discovery and development process. Furthermore, the number of clinical trials conducted each year continues to rise, with their data dynamically evolving under the influence of various external factors, such as technological innovations and global events like the COVID-19 pandemic. This ever-changing landscape underscores the need for updated trial outcome labels to develop robust and high-performing clinical trial outcome prediction models.
@@
Trial Outcome Definition: Clinical trial outcomes are multifaceted and have diverse implications. These outcomes can involve meeting the primary endpoint as defined in the study, advancing to the next phase of the trial, obtaining regulatory approval, impacting the financial outcome for the sponsor (either positively or negatively), and influencing patient outcomes such as adverse events and trial dropouts. Our paper follows the previous conventions [11, 12, 3, 13, 14, 15] and defines the trial outcome as a binary indicator, showing whether the trial achieves its primary endpoints and can progress to the next stage of drug development. For example, for Phase 1 and 2 trials, success may mean moving to the next phase, such as from Phase 1 to Phase 2, and from Phase 2 to Phase 3. In Phase 3, success is measured by regulatory approval.
@@
Recent efforts have been made to predict specific aspects of clinical trial outcomes. These include employing electroencephalographic (EEG) measures to predict the effects of antidepressant treatments [16], optimizing drug toxicity predictions based on drug and target properties [17], and using phase II trial results to anticipate phase III trial outcomes [18]. Additionally, there is a growing interest in developing comprehensive methods for predicting trial outcomes. For example, predicting drug approvals for 15 disease groups by analyzing drug and clinical trial features using classical machine learning techniques [14], using multimodal drug structure and text information to predict outcomes based on a supervised set of data [11, 12], an algorithm that computes the probability of technical success via asking experts a standardized questionnaire [19], and multimodal trial outcome prediction via omics, text, clinical trial design, and small molecule properties [15].
@@
To address these limitations, we propose CTO, the first large-scale, publicly available, open-source, and fully reproducible dataset of clinical trial outcomes. We compiled a comprehensive trial knowledge base by integrating diverse data sources, including trial-related publications, news headlines, stock prices of trial sponsors, and other relevant trial metrics. From this knowledge base, we derive weakly supervised trial outcome labels using the CTO’s automated labeling framework, which aggregates labels from trial phase linkages, LLM interpretations on publications, sentiment analysis of news headlines, p-values, and trial status, among other indicators. We validate CTO’s labeling accuracy against previously published trial outcomes, achieving an F1 score of 91 across all trial phases, closely aligned with human-annotated clinical trial outcomes. Additionally, aided by the CTO’s framework, we manually curated a set of approximately 10,000 recent trials completed between 2020 and 2024. To ensure reproducibility and keep pace with rapidly evolving trial data, we will share regular updates of our trial knowledge base and labels generated by CTO’s automated labeling framework.
Figure 1: Overview of our CTO benchmark pipeline. (A) Trial selection criteria used to curate a targeted dataset from the CTTI database, narrowing from an initial 513.4K trials to 125.8K trials in CTO to a final manually curated set of 11K trials. (B) Trial knowledge base creation pipeline, including extraction of linked PubMed abstracts, recent abstract searches, trial metrics, news articles, and stock prices. (C) Clinical trial outcome labeling process, integrating the CTO automated labeling framework assisted manual labeling to generate curated gold standard labels. (D) The manual curation process, our rule-based annotation criteria, followed by manual checking on challenging clinical trials.
@@ Methods
Our study focused primarily on clinical trials involving drugs and biologics. Starting with a dataset of approximately 513,485 clinical trials from the Clinical Trial Transformation Initiative (CTTI) database [20], we applied a series of selection criteria to create a refined subset as illustrated in Figure 1A. First, we selected only trials with interventions categorized as drugs or biologics, reducing the dataset to 209,124 trials. Next, we excluded active trials and those still recruiting to focus exclusively on completed studies, resulting in 145,464 trials. Subsequently, trials without phase information were removed, further narrowing the dataset to 125,840 trials, representing the number of trials labeled using our CTO’s automated labeling framework. To construct a manually curated test set, we then selected trials completed after 2020, yielding a subset of approximately 30,900 recent trials. Importantly, we ensured this set did not overlap with the TOP benchmark[11]. Finally, we sampled different trial phases and conditions, producing a manually curated set of 11,012 trials.
@@
Based on the CTTI data, we generate a trial knowledge base by aggregating and linking various data sources to the trials, as shown in Figure 1B. Our pipeline enables regular updates of the knowledge base given the updated CTTI data on a monthly frequency. The trial knowledge base extraction pipeline consists of four components: 1) PubMed abstract extraction, 2) News extractions, 3) Trial metrics, and 4) Stock price information extraction.
Pubmed abstract extraction: PubMed abstracts have been automatically linked to trials by the Clinical Trials Transformation Initiative (CTTI)[21, 22, 23] as well as other efforts [24]. Despite these sources, we found that many recent abstracts have not linked to their corresponding trials. To alleviate the linking issues, we perform the following procedure. First, we extracted all linked PubMed abstracts for each trial through the National Center for Biotechnology Information (NCBI) API  [^3] https://ncbi.nlm.nih.gov/, and the statistics of the linked abstracts are given in the Supplementary. These abstracts can be categorized into 1) Background, 2) Derived, and 3) Results. Since we are interested in clinical trial outcomes, we only utilized abstracts in the Derived and Results categories. To enrich the knowledge base further, we performed regular searches on NCBI using NCT IDs to find newly published abstracts, linking any relevant ones to their associated trials. The initial count of drug or biologic trials with extracted PubMed abstracts linked from CTTI data was 49,277. Our additional search and extraction process linked 10,232 new trials with their recent PubMed abstracts. Ultimately, our knowledge base comprises 41,228 Results-type abstracts, 65,327 Derived-type abstracts, and 20,512 newly linked abstracts.
News extractions: In our initial analysis, we identified news sources as a valuable resource for determining the outcome of clinical trials, especially since most trials in the CTTI dataset lack reported results. We extracted news articles related to completed drug and biologics trials by web scraping and linked them to the trials in our knowledge base. We sent requests to Google News for all completed drug and biologic trials (excluding phase 4), which accounted for around 85K trials. Due to rate limitations, we used an external API (SerpAPI) to streamline this process. Our findings revealed that most trials had no corresponding news coverage; only 3,688 were associated with news articles.
Trial metrics: We extracted trial metrics from CTTI[20] and mining clinicaltrials.gov. Specifically, we extracted nine numerical metrics–the number of patients, sponsors, sites, amendments, patient drops, deaths, serious adverse events, all adverse events, and the number of days the trial has been last updated after completion. We also extracted three binary metrics–whether results were reported, the completion/terminated status of the trial, and finally, whether the trial had a significant p-value. Detailed descriptions of each are in the supplementary.
@@
We developed a clinical trial outcome labeling pipeline that first assigns rule-based labels. In this paper, labels or labeling refer to the machine learning concept related to the ground-truth prediction target, unrelated to drug labels. First, trials with statuses indicating ‘terminated’, ‘withdrawn’, ‘suspended’, ‘no longer available’} were labeled as failures. Trials labeled as ‘approved for marketing’ were marked as successes. For trials with ‘completed’ as status, we examined the reported results: if any reported p-value for a primary endpoint was $<$0.05, the trial was labeled as a success. Unlabeled trials are then automatically labeled. The main idea behind our automated labeling framework is to combine multiple sources of labels with varying levels of accuracy. We call these sources of labels “weakly supervised labeling functions” or LFs for short[25]. Following this, we apply a label model that leverages our sources to refine labels and produce denoised pseudo-labels. Each LF, such as detecting efficacy based on p-value thresholds, assigns labels or abstains when uncertain. Certain metrics, such as adverse events, are also subject to phase-specific thresholds to account for variations in participant numbers across trial phases.
@@ Results
Table 1: Evaluation of automatically aggregated knowledge base data sources. Agreement of CTO with TOP (on TOP test split). $CTO_{MV}$ denotes Majority Vote, $CTO_{DP}$ denotes Data Programming [25], and $CTO_{RF}$ denotes Random Forest. Note that RF actively trained on the training human labels, whereas MV and DP are not. Bolded denotes best performance.
@@
I $CTO_{MV}$ 0.726 0.490
$CTO_{DP}$ 0.870 0.700
$CTO_{RF}$ 0.913 0.790
II $CTO_{MV}$ 0.689 0.430
$CTO_{DP}$ 0.856 0.623
$CTO_{RF}$ 0.878 0.693
III $CTO_{MV}$ 0.904 0.606
$CTO_{DP}$ 0.921 0.582
$CTO_{RF}$ 0.941 0.710
All $CTO_{MV}$ 0.793 0.529
$CTO_{DP}$ 0.884 0.646
$CTO_{RF}$ 0.909 0.729
We evaluate our knowledge base and labeling framework by comparing the agreement between our automatically aggregated labels against the expert-annotated TOP test set. Table 1 summarizes the performance of various label aggregation methods across different trial phases. The label aggregation methods evaluated include Majority Vote (CTOMV), Data Programming (CTODP), and Random Forest-based aggregation (CTORF). Across all trial phases, the Random Forest-based approach consistently outperforms the other methods, achieving the highest scores in both F1 and Cohen’s kappa metrics. CTORF achieves an F1 score of $0.909$ and a Cohen’s kappa ($\kappa$) of $0.729$ when evaluated across all phases, indicating substantial agreement [26]. Also, the phase-specific analysis further underscores the robustness of CTORF. It achieves F1 scores of $0.913$, $0.878$, and $0.941$ in phases I, II, and III, respectively. The F1 score of 0.909 is much higher than previous trial outcome prediction models [27, 12].
@@
Automated labeling is not a substitute for human expertise, so we manually annotated over 2,500 challenging cases using our knowledge base to create a high-confidence gold standard set for evaluation. The process for developing this benchmark is illustrated in Figure 1D. We applied the same rule-based termination criteria and p-value threshold of less than 0.05 as used in the automated labeling step to filter out the easier cases. For the remaining unlabeled trials, we engaged three clinical trial experts to review relevant publications, news articles, and other reported outcomes to accurately determine the trial outcomes. In this benchmark evaluation, we compared the labels generated by the CTO automated labeling framework against the human-generated labels. The CTO framework demonstrated strong overall performance, achieving an F1 score of 97.1. Notably, even in challenging cases (after applying p-value filtering and status checks), where accurate labeling is more complex, the CTO framework maintained an F1 score of 72.7.
@@
Consistent High Prediction Performance Over Time
@@
Figure 3: (A) Year-by-year comparison of trial outcome prediction performance for baseline models trained on the TOP dataset, the constantly updated CTO dataset, and our manually curated set, evaluated on trials completed from 2021 to 2024. The TOP dataset is fully utilized for training, whereas for the CTO and manually curated sets, only trials completed before each test year are used for training. (B) Top 10 Condition-wise performance of baselines on manual labels, with models trained on data $<$ 2022 and tested on $\geq$ 2022 using TOP, CTO, and manually curated set. (C) Similar to B, Condition-wise performance on the next 10 (Top 11-20 conditions)
@@
Given that we have manual labels for recent trials, we conducted experiments comparing the impact of our constantly updated CTO dataset against the previously publicly available fixed benchmark TOP dataset. We trained standard baseline trial outcome prediction models on the entire TOP dataset (train + valid + test). We assessed their year-by-year performance on trials completed between 2021 and 2024 in our manually curated dataset. For the CTO dataset, the baselines were trained using only trials completed before each test year, with training labels generated through our automated labeling framework. Similarly, we trained the baselines on the manually curated dataset using human-labeled outcomes for trials completed before each test year. The phase-wise performance comparison of these baseline models across the different training datasets is presented in Figure 3A.
@@
Baseline models trained using labels from the CTO dataset consistently outperform their counterparts trained on the TOP dataset across all years and trial phases. The CTO dataset significantly enhances model performance compared to the TOP dataset, underscoring the importance of constantly updated labels. This improvement is mainly attributable to distribution shifts caused by evolving trial practices, therapeutic areas, and advancements in trial designs over time.
@@
CTO provides a scalable solution for labeling recent trial data, overcoming the infeasibility and time demands of obtaining new human labels. By achieving high agreement with human labels, the CTO enables models to obtain close-to-human-level performance by leveraging large datasets to adapt to the evolving landscape of clinical trials.
Performance of Condition-Specific Outcome Predictions
@@
Figure 3B and C present the performance of baseline models trained using our human-labeled data across the top 20 conditions in the dataset (3B shows conditions 1-10, and 3C shows conditions 11-20). Here, we train the baselines on trials completed before 2022 over TOP, CTO, and manually curated. The baselines are tested on the manually curated trials completed after 2022. The performance of the baseline models trained on CTO closely follows the models trained on our human labels. The performance demonstrates that the labels produced by CTO align closely with manual annotations despite being generated independently through an automated process. Moreover, these CTO-based models consistently outperform those trained on the TOP dataset, demonstrating the value of using a more current dataset. Furthermore, CTO facilitates the development of condition-specific clinical trial outcome prediction models by rapidly generating automated labels for condition-specific datasets, significantly reducing the time and effort involved in manual labeling, with CTO-trained baselines matching or even exceeding baselines trained on manually-curated in some cases.